# Gradients of Functions of Large Matrices

### v1

@@ 1 Introduction
Evaluating Equation 1 is crucial for building large machine learning models, e.g., Bayesian neural nets: A common hyperparameter-calibration loss of a (Laplace-approximated) Bayesian neural network involves the log-determinant of the generalised Gauss–Newton matrix [13]
@@ 4 The method: Adjoints of the Lanczos and Arnoldi iterations
In other words, we “linearise” the constraint $\mathbf{c}(\theta,x)=0$. The adjoint method [78] proceeds by “transposing” this linearisation as follows. Let $\rho$ be a loss that depends on $y$ with gradient $\nabla_{y}\rho$. Then, for all Lagrange multipliers $\lambda$ with the same shape as the outputs of $\mathbf{c}(\cdot)$, we know
$\mathrm{d}\rho=\langle\nabla_{x}\rho,\mathrm{d}x\rangle+\langle\lambda,\mathrm{d}\mathbf{c}\rangle=\langle\nabla_{x}\rho+(D_{x}\mathbf{c})^{\top}\lambda,\mathrm{d}x\rangle+\langle(D_{\theta}\mathbf{c})^{\top}\lambda,\mathrm{d}\theta\rangle.$ (9)
By matching Equation 9 to $\mathrm{d}\rho=\langle\nabla_{\theta}\rho,\mathrm{d}\theta\rangle$, we conclude that if $\lambda$ solves the adjoint system
@@ 4.2 Matrix-free implementation
To compute $\nabla_{A}\rho$ and $\nabla_{v}\rho$, we need to solve the adjoint systems. When comparing the forward constraints to the adjoint systems, similarities emerge: for instance, the adjoint system of the Arnoldi iteration follows the same $A^{(\top)}H-\Lambda H^{(\top)}+\text{rest}=0$ structure as the forward constraint. This structure suggests deriving a recursion for the backward pass that mirrors that of the forward pass. Appendix E contains this derivation and contrasts the resulting algorithm with that of the forward pass. The main observation is that the complexity of the adjoint passes for Lanczos and Arnoldi mirrors that of the forward passes. Gradients can be implemented purely with matrix-vector products, which explains the significant performance gains observed in Figure 3.
@@
$\nabla_{\theta}\rho=\sum_{k=1}^{K}\nabla\left[\theta\mapsto(e_{k})^{\top}Q^{\top}A(\theta)^{\top}\Lambda e_{k}\right]$ (19)
@@
Table 2: Accuracy loss when differentiating the Arnoldi iteration on a Hilbert matrix in double precision (Appendix F).
@@
The main takeaway from Sections 4.1 and 4.2 is that now, we do not only have closed-form expressions for the gradients of Arnoldi and Lanczos iterations (Theorems 4.1 and 4.2), but that we can compute them in the same complexity as the forward pass, in a numerically stable way, and evaluate parameter-gradients in linear time- and space-complexity (Corollary 4.3). While some of the derivations are somewhat technical, the overall approach follows the general template for the adjoint method relatively closely. The resulting algorithm beats backpropagation “through” the iterations by a margin in terms of speed (Figure 3) and enjoys the same stability gains from reorthogonalisation as the forward pass (Table 2). Our open-source implementation of reverse-mode differentiable Lanczos and Arnoldi iterations can be installed via pip install [redacted]. Next, we put this code to the test on three challenging machine-learning problems centred around functions of matrices to see how it fares against state-of-the-art differentiable implementations of exact Gaussian processes (Section 5), differential equation solvers (Section 6), and Bayesian neural networks (Section 7).
@@ 6 Case study: Physics-informed machine learning with PDEs
Figure 5: Arnoldi’s superior convergence on the forward pass (A1) is inherited by the gradients (A2) and ultimately leads to fast training (B). For training, Arnoldi uses ten matrix-vector products, and the other two use 15 matrix-vector products (so they have equal error $\approx 10^{-4}$ in A1 and A2.)
@@ 7 Case study: Calibrating Bayesian neural networks
AUROC (out-of-dist.) $\uparrow$ 0.9 $\pm$ 0.03 0.9 $\pm$ 0.03 0.5 $\pm$ 0.01
@@
This work was supported by a research grant (42062) from VILLUM FONDEN. The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606)

### v2

@@ 1 Introduction
Evaluating Equation 1 is crucial for building large machine learning models, e.g., Bayesian neural networks: A common hyperparameter-calibration loss of a (Laplace-approximated) Bayesian neural network involves the log-determinant of the generalised Gauss–Newton matrix [13]
@@ 4 The method: Adjoints of the Lanczos and Arnoldi iterations
In other words, we “linearise” the constraint $\mathbf{c}(\theta,x)=0$. The adjoint method [78] proceeds by “transposing” this linearisation as follows. Let $\rho$ be a loss that depends on $y$ with gradient $\nabla_{y}\rho$ and recall the gradient identity from the “Notation” paragraph above. Then, for all Lagrange multipliers $\lambda$ with the same shape as the outputs of $\mathbf{c}(\cdot)$, we know that since $\mathbf{c}=0$ implies $\mathrm{d}\mathbf{c}=0$,
$\mathrm{d}\rho=\langle\nabla_{x}\rho,\mathrm{d}x\rangle=\langle\nabla_{x}\rho,\mathrm{d}x\rangle+\langle\lambda,\mathrm{d}\mathbf{c}\rangle=\langle\nabla_{x}\rho+(D_{x}\mathbf{c})^{\top}\lambda,\mathrm{d}x\rangle+\langle(D_{\theta}\mathbf{c})^{\top}\lambda,\mathrm{d}\theta\rangle$ (9)
must hold. By matching Equation 9 to $\mathrm{d}\rho=\langle\nabla_{\theta}\rho,\mathrm{d}\theta\rangle$ (this time, regarding $\rho$ as a function of $\theta$, not of $x$; recall the “Notation” paragraph), we conclude that if $\lambda$ solves the adjoint system
@@ 4.2 Matrix-free implementation
To compute $\nabla_{A}\rho$ and $\nabla_{v}\rho$, we need to solve the adjoint systems. When comparing the forward constraints to the adjoint systems, similarities emerge: for instance, the adjoint system of the Arnoldi iteration follows the same $A^{(\top)}H-\Lambda H^{(\top)}+\text{rest}=0$ structure as the forward constraint. This structure suggests deriving a recursion for the backward pass that mirrors that of the forward pass. Appendix E contains this derivation and contrasts the resulting algorithm with that of the forward pass. The main observation is that the complexity of the adjoint passes for Lanczos and Arnoldi mirrors that of the forward passes. Gradients can be implemented purely with matrix-vector products, which is helpful because it makes our custom backward pass as matrix-free as backpropagation “through” the forward pass would be. This matrix-free implementation in combination with the efficient recursions in Theorems 4.1 and 4.2 explains the significant performance gains of our method compared to naive backpropagation, observed in Figure 3.
@@
$\nabla_{\theta}\rho=\sum_{k=1}^{K}\nabla\left[\theta\mapsto(e_{k})^{\top}Q^{\top}A(\theta)^{\top}\Lambda e_{k}\right],$ (19)
@@
Table 2: Accuracy loss when differentiating the Arnoldi iteration on a Hilbert matrix in double precision (ϕ::italic-$\phi:$ decompose with a full-rank Arnoldi iteration, then reconstruct the original matrix; measure $\|\partial\phi-I\|$; details in Appendix F).
@@
The main takeaway from Sections 4.1 and 4.2 is that now, we do not only have closed-form expressions for the gradients of Arnoldi and Lanczos iterations (Theorems 4.1 and 4.2), but that we can compute them in the same complexity as the forward pass, in a numerically stable way, and evaluate parameter-gradients in linear time- and space-complexity (Corollary 4.3). While some of the derivations are somewhat technical, the overall approach follows the general template for the adjoint method relatively closely. The resulting algorithm beats backpropagation “through” the iterations by a margin in terms of speed (Figure 3) and enjoys the same stability gains from reorthogonalisation as the forward pass (Table 2). Our open-source implementation of reverse-mode differentiable Lanczos and Arnoldi iterations can be installed via “pip install matfree”. Next, we put this code to the test on three challenging machine-learning problems centred around functions of matrices to see how it fares against state-of-the-art differentiable implementations of exact Gaussian processes (Section 5), differential equation solvers (Section 6), and Bayesian neural networks (Section 7).
@@ 6 Case study: Physics-informed machine learning with PDEs
Figure 5: Arnoldi’s superior convergence on the forward pass (A1) is inherited by the gradients (A2; mind the shared $y$-axis) and ultimately leads to fast training (B). For training, Arnoldi uses ten matrix-vector products, and the other two use 15 (so they have equal error $\approx 10^{-4}$ in A1 and A2.)
@@ 7 Case study: Calibrating Bayesian neural networks
AUROC (out-of-dist.) $\uparrow$ 0.9 $\pm$ 0.03 0.9 $\pm$ 0.03 0.5 $\pm$ 0.010
@@
This work was supported by a research grant (42062) from VILLUM FONDEN. The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606). This project received funding from the European Research Council (ERC) under the European Union’s Horizon programme (grant agreement 101125993).